# Bayesian Group Regularization in Generalized Linear Models with a Continuous Spike-and-Slab Prior thanks: Keywords and phrases: generalized linear model, group sparsity, spike-and-slab, spike-and-slab group lasso, posterior contraction, rate of convergence, variable selection

### v8

@@ 1.1 Motivation
Nowadays, it is common to collect datasets where the total number of covariates $p$ is large. In this “large $p$” setting, the covariates also often exhibit a grouping structure. For example, in genomic data, genes within the same biological pathway function together as a group to affect a clinical phenotype such as disease status or survival time [18]. At the individual gene level, mutations in an amino acid sequence can also be represented with group structure [34]. In our motivating data application in Section 7, each position of a protease gene sequence is represented by a group of binary variables, where each binary variable indicates the presence or absence of a specific amino acid. In MRI imaging, voxels within the same brain region naturally form a group [24, 46]. In semiparametric GLMs, continuous functions of the covariates are often estimated by groups of nonlinear basis functions [26]. Finally, categorical covariates with multiple levels can be represented as groups of dummy variables for each non-baseline category [9, 29].
@@ 1.3 Notation and preliminaries
$-\nabla^{2}\ell_{n}(\bm{\beta})=\mathbf{X}^{\top}\bm{\Sigma}(\bm{\beta})\mathbf{X},$ (1.6)
@@ 2.2 Bayesian estimation
[…]
@@
[…]
Local MAP estimators (or local posterior modes) are useful point estimates to consider because standard optimization algorithms can be used to rapidly obtain an estimate $\widehat{\bm{\beta}}$. These optimization algorithms are often faster and much more scalable than MCMC algorithms. In general, local MAP estimators may not be unique, and the likelihood function $\ell_{n}(\bm{\beta})$ may be unbounded. However, if the prior $\pi(\bm{\beta})$ in (2.5) is proper, does not depend on the observed data, and sufficiently restricts the parameter space for $\bm{\beta}$ (e.g. by shrinking most of the coefficients in $\bm{\beta}$ towards zero), then the posterior density $\pi(\bm{\beta}\mid\mathbf{Y})$ will be proper and bounded, i.e. ${\color[rgb]{0,0,0}\sup_{\bm{\beta}}}\pi(\bm{\beta}\mid\mathbf{Y})<\infty$. This enables us to find a local MAP estimator (2.6).
@@ 3.1 Convergence rate of the MAP estimator
By the continuous mapping theorem, $f(\widehat{\bm{\beta}})$ is also a consistent estimator of $f(\bm{\beta}_{0})$ for any function $f$ that is continuous at $\bm{\beta}_{0}$. For example, let $S\subset\{1,\ldots,p\}$, and let $\bm{\beta}_{S}$ be the subvector of $\bm{\beta}$ with indices in $S$. Taking $f(\bm{\beta})=\mathbf{A}_{S}\bm{\beta}$, where $\mathbf{A}_{S}\in\mathbb{R}^{|S|\times p}$ is a matrix with unit vector rows such that $\mathbf{A}_{S}\bm{\beta}=\bm{\beta}_{S}$, it follows that $\widehat{\bm{\beta}}_{S}$ is a consistent estimator of $\bm{\beta}_{S}$.
@@ 5.1 EM algorithm for MAP estimation
We use a variant of the EMVS algorithm [36] to iteratively solve for the MAP estimator $(\widehat{\bm{\beta}},\widehat{\theta})$ in the optimization problem,
@@ 5.3 Gibbs sampling for fully Bayesian inference
[…]
@@
~$\widetilde{\mathcal{L}}_{i}(\bm{\beta})\approx\frac{\left\{\exp(\mathbf{x}_{i}^{\top}\bm{\beta}-\log M)\right\}^{0.5y_{i}}}{\left\{1+\exp(\mathbf{x}_{i}^{\top}\bm{\beta}-\log M)\right\}^{M}},$ (5.11)
@@ 6.1 Setup and performance metrics
[…]
@@ 6.4 Fully Bayesian inference
    the coverage probability (CP) of the 95% posterior credible intervals (CIs) for the 40 true non-null coefficients in $\bm{\beta}$, i.e. the proportion of CIs for the nonzero regression coefficients $\{\beta_{j}:\beta_{j}\neq 0\}$ that contained the true $\beta_{j}$. These intervals were constructed using the 0.025 and 0.975 sample quantiles $[q_{0.025}(\beta_{j}),q_{0.975}(\beta_{j})]$ of the saved 2000 MCMC samples for each $\beta_{j}$.
@@ Appendix A Proofs for Section 3
=$\overset{\Delta}{=}(\textrm{I})+(\textrm{II})<0,$ (A.10)
@@
[…]
=$\overset{\Delta}{=}J_{1}+J_{2},$ (A.11)
@@
[…]
@@
where $\qquad\qquad\textrm{where }\widetilde{\bm{\beta}}_{g}\textrm{ lies on line segment between }\bm{\beta}_{0g}\textrm{ and }\bm{\beta}_{0g}+\epsilon_{n}\bm{\Delta}_{g},$
@@
[…]
@@
$\lVert\mathbf{X}_{g}^{\top}\textrm{diag}(\xi^{\prime}(\mathbf{X}\widehat{\bm{\beta}}))[\mathbf{Y}-{\color[rgb]{0,0,0}b^{\prime}(\xi({\color[rgb]{0,0,0}\mathbf{X}\widehat{\bm{\beta}}}))]}\rVert_{2}<\lambda_{\bm{\theta}}^{\star}(\bm{0}_{m_{g}})~{}~{}\textrm{for all }g\in S_{0}^{c}.$ (A.18)
@@
Note that $\lambda_{\max}(\lVert\bm{\widehat{\beta}}_{g}\rVert_{2}^{-2}\bm{\widehat{\beta}}_{g}\bm{\widehat{\beta}}_{g}^{\top})=1$, since $\bm{\widehat{\beta}}_{g}\bm{\widehat{\beta}}_{g}^{\top}$ has the same nonzero eigenvalue as $\bm{\widehat{\beta}}_{g}^{\top}\bm{\widehat{\beta}}_{g}=\lVert\bm{\widehat{\beta}}_{g}\rVert_{2}^{2}$. Meanwhile, the only eigenvalue of $\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{-1}\mathbf{I}_{m_{g}}$ is $\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{-1}$, and the only eigenvalues of $\lVert\bm{\widehat{\beta}}_{g}\rVert_{2}^{-3}\bm{\widehat{\beta}}_{g}\bm{\widehat{\beta}}_{g}^{\top}$ are 0 and $\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{-1}$. Consequently, $-\lambda_{\min}(\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{-1}\mathbf{I}_{m_{g}}-\lVert\bm{\widehat{\beta}}_{g}\rVert_{2}^{-3}\bm{\widehat{\beta}}_{g}\bm{\widehat{\beta}}_{g}^{\top})\leq 0$. Using Weyl’s inequality and the fact that $\lambda_{\max}(-\mathbf{A})=-\lambda_{\min}(\mathbf{A})$ for a square matrix $\mathbf{A}$, it follows from (A) that
@@
[…]
@@
$P\left(\exists~{}g\in S_{0}^{c},~{}~{}\bigg{\lVert}\frac{\partial\ell_{n}(\widehat{\bm{\beta}})}{\partial\bm{\beta}_{g}}\bigg{\rVert}_{2}\geq\lambda_{g}^{\star}(\bm{0}_{m_{g}})\right)\rightarrow 0.$
@@
$P\left(\exists~{}g\in S_{0}^{c},~{}~{}\lVert h_{g}(\widehat{\bm{\beta}})\rVert_{2}\geq\lambda_{g}^{\star}(\bm{0}_{m_{g}})\right)$
$\qquad\leq P\left(\exists~{}g\in S_{0}^{c},~{}~{}\lVert h_{g}(\bm{\beta}_{0})\rVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)+P\left(\exists~{}g\in S_{0}^{c},~{}~{}\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\rVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)$
$\qquad\leq\sum_{g=s_{0}+1}^{G}P\left(\lVert h_{g}(\bm{\beta}_{0})\rVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)+\sum_{g=s_{0}+1}^{G}P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\rVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)$
=$\qquad\overset{\Delta}{=}T_{1}+T_{2}.$ (A.21)
@@
$\leq 2\exp\left(-\frac{\lambda_{0}^{2}/4D^{2}m_{\max}}{2nG+L\lambda_{0}/D{\color[rgb]{0,0,0}\log G}\sqrt{m_{\max}}}\right)$
$=2\exp\left(-\frac{\lambda_{0}^{2}}{8D^{2}m_{\max}nG+4L\lambda_{0}D\sqrt{m_{\max}}/{\color[rgb]{0,0,0}\log G}}\right)$ (A.23)
@@
$T_{1}<2\exp\left(\log G-\frac{\lambda_{0}^{2}}{8D^{2}m_{\max}nG+4L\lambda_{0}D\sqrt{m_{\max}}{\color[rgb]{0,0,0}/\log G}}\right)\rightarrow 0,$ (A.24)
@@
We next examine each of the summands in $T_{2}$ in (A). In light of (A), we have that for each $g\in S_{0}^{c}$ and sufficiently large $n$,
$P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\lVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)$ $<P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\lVert_{2}\geq\frac{\lambda_{0}}{2}\right)$
$\leq P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})\rVert_{2,\infty}\lVert\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}\geq\frac{\lambda_{0}}{2}\right)$
$\leq P\left(\lVert\mathbf{X}_{S_{0}}^{\top}\bm{\Sigma}(\widetilde{\bm{\beta}})\mathbf{X}_{g}\rVert_{2,\infty}\lVert\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}\geq\frac{\lambda_{0}}{2}\right)$
$\lesssim\frac{2n}{\lambda_{0}},$ (A.25)
To arrive at (A), we used the fact that $\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})=\mathbf{X}_{S_{0}}^{\top}\bm{\Sigma}(\widetilde{\bm{\beta}})\mathbf{X}_{g}$ and Assumption (A3)(iii) that for all $g\in\mathcal{S}_{0}^{c}$, $\lVert\mathbf{X}_{S_{0}}^{\top}\bm{\Sigma}(\widetilde{\bm{\beta}})\mathbf{X}_{g}\rVert_{2,\infty}=O(n)$. From (A) and the fact that $\lambda_{0}\asymp G^{c},c>2$, and $G\gg n$, we have as an upper bound for $T_{2}$ in (A),
$T_{2}\prec\frac{2Gn}{\lambda_{0}}\asymp\frac{Gn}{G^{c}}=\frac{n}{G^{c-1}}\rightarrow 0\textrm{ as }n\rightarrow\infty.$ (A.26)

### v9

@@ 1.1 Motivation
Nowadays, it is common to collect datasets where the total number of covariates $p$ is large. In this “large pp” setting, the covariates also often exhibit a grouping structure. For example, in genomic data, genes within the same biological pathway function together as a group to affect a clinical phenotype such as disease status or survival time [18]. At the individual gene level, mutations in an amino acid sequence can also be represented with group structure [34]. In our motivating data application in Section 7, each position of a protease gene sequence is represented by a group of binary variables, where each binary variable indicates the presence or absence of a specific amino acid. In MRI imaging, voxels within the same brain region naturally form a group [24, 46]. In semiparametric GLMs, continuous functions of the covariates are often estimated by groups of nonlinear basis functions [26]. Finally, categorical covariates with multiple levels can be represented as groups of dummy variables for each non-baseline category [9, 29].
@@ 1.3 Notation and preliminaries
$\nabla^{2}\ell_{n}(\bm{\beta})=\mathbf{X}^{\top}\bm{\Sigma}(\bm{\beta})\mathbf{X},$ (1.6)
@@ 2.2 Bayesian estimation
$\pi(\bm{\beta}\mid\mathbf{Y})=\frac{\exp(\ell_{n}(\bm{\beta}))\pi(\bm{\beta})}{\int\exp(\ell_{n}(\bm{\beta}))\pi(\bm{\beta})d\bm{\beta}},$ (2.4)
@@
$\widehat{\bm{\beta}}\text{ such that }\nabla\log\pi(\widehat{\bm{\beta}}\mid\mathbf{Y})=\bm{0}_{p}.$ (2.6)
Local MAP estimators (or local posterior modes) are useful point estimates to consider because standard optimization algorithms can be used to rapidly obtain an estimate $\widehat{\bm{\beta}}$. These optimization algorithms are often faster and much more scalable than MCMC algorithms. In general, local MAP estimators may not be unique, and the log-likelihood function $\ell_{n}(\bm{\beta})$ may be unbounded. However, if the prior $\pi(\bm{\beta})$ in (2.5) is proper, does not depend on the observed data, and sufficiently restricts the parameter space for $\bm{\beta}$ (e.g. by shrinking most of the coefficients in $\bm{\beta}$ towards zero), then the posterior density $\pi(\bm{\beta}\mid\mathbf{Y})$ will be proper and bounded, i.e. ${\color[rgb]{0,0,0}\sup_{\bm{\beta}}}\pi(\bm{\beta}\mid\mathbf{Y})<\infty$. This enables us to find a local MAP estimator (2.6).
@@ 3.1 Convergence rate of the MAP estimator
[…]
@@ 5.1 EM algorithm for MAP estimation
[…]
@@ 5.3 Gibbs sampling for fully Bayesian inference
$\frac{(e^{\psi})^{a}}{(1+e^{\psi})^{b}}=2^{-b}e^{\kappa\psi}\int_{0}^{\infty}e^{-\omega\psi^{2}/2}p(\omega)d\omega,$ (5.8)
@@
$\widetilde{\mathcal{L}}_{i}(\bm{\beta})\approx\frac{\left\{\exp(\mathbf{x}_{i}^{\top}\bm{\beta}-\log M)\right\}^{0.5y_{i}}}{\left\{1+\exp(\mathbf{x}_{i}^{\top}\bm{\beta}-\log M)\right\}^{M}},$ (5.11)
@@ 6.1 Setup and performance metrics
In the simulated genetic association studies (Experiment 4 in Sections 6.2 and 6.3), the responses are either binary or count phenotypes, and the covariates are simulated SNPs. SNPs are categorical variables coded as one of three values $\{$“0”, “1”, or “2”}\}, depending on the number of minor alleles present [9]. We thus represent each SNP as a factor with two levels, i.e. a group of two indicator variables. Assuming that “2” is the baseline, we can represent each $j$th SNP $x_{j}$ with two dummy variables $\mathbb{I}(x_{j}=0)$ and $\mathbb{I}(x_{j}=1)$. If $x_{j}=2$, then $\mathbb{I}(x_{j}=0)=\mathbb{I}(x_{j}=1)=0$.
@@ 6.4 Fully Bayesian inference
    the coverage probability (CP) of the 95% posterior credible intervals (CIs) for the 40 true non-null coefficients in $\bm{\beta}$, i.e. the proportion of CIs for the nonzero regression coefficients $\{\beta_{j}:\beta_{j}\neq 0\}$ that contained the true $\beta_{j}$. These intervals were constructed using the 0.025 and 0.975 sample quantiles $q_{0.025}(\beta_{j}),q_{0.975}(\beta_{j})]$ of the saved 2000 MCMC samples for each $\beta_{j}$.
@@ Appendix A Proofs for Section 3
$\overset{\Delta}{=}(\textrm{I})+(\textrm{II})<0,$ (A.10)
@@
$(\textrm{I})$ $=\epsilon_{n}\nabla_{S_{0}}\ell_{n}(\bm{\beta}_{0})^{\top}\bm{\Delta}_{S_{0}}+\frac{\epsilon_{n}^{2}}{2}\bm{\Delta}_{S_{0}}^{\top}\nabla^{2}_{S_{0}}{\color[rgb]{0,0,0}\ell_{n}}(\widetilde{\bm{\beta}})\bm{\Delta}_{S_{0}}$
$\overset{\Delta}{=}J_{1}+J_{2},$ (A.11)
@@
Now we turn our attention to upper-bounding $(\textrm{II})$ in (A). We first focus on bounding each of the summands in (II) from above. We first rewrite $\textrm{pen}(\bm{\beta}_{g})$ in (A.1) as
@@
[…]
@@
and so we can further bound (A) from above by $\epsilon_{n}(1+\lambda_{1})\lVert\bm{\Delta}_{g}\rVert_{2}$. Therefore, an upper bound for $(\textrm{II})$ in (A) is
@@
$\lVert\mathbf{X}_{g}^{\top}\textrm{diag}(\xi^{\prime}(\mathbf{X}\widehat{\bm{\beta}}))[\mathbf{Y}-{\color[rgb]{0,0,0}b^{\prime}(\xi({\color[rgb]{0,0,0}\mathbf{X}\widehat{\bm{\beta}}}))]}\rVert_{2}<\lambda_{\theta}^{\star}(\bm{0}_{m_{g}})~~\textrm{for all }g\in S_{0}^{c}.$ (A.18)
@@
Note that $\lambda_{\min}(\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{-1}\mathbf{I}_{m_{g}}-\lVert\bm{\widehat{\beta}}_{g}\rVert_{2}^{-3}\bm{\widehat{\beta}}_{g}\bm{\widehat{\beta}}_{g}^{\top})=0$, since $\mathbf{I}_{m_{g}}-\rVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{-2}\widehat{\bm{\beta}}_{g}\widehat{\bm{\beta}}_{g}^{\top}$ is the annihilator matrix. Using the fact that $\lambda_{\max}(-\mathbf{A})=-\lambda_{\min}(\mathbf{A})$ for a square matrix $\mathbf{A}$, it follows from (A) that
@@
$\lesssim-n+\frac{c(\widehat{\bm{\beta}}_{g})(\lambda_{0}-\lambda_{1})^{2}}{(1+c(\widehat{\bm{\beta}}_{g}))^{2}},~~\text{where }c(\bm{\beta}_{g})\text{ is as in }(\text{A}.15)$
@@
$P\left(\exists~g\in S_{0}^{c},~~\bigg{\lVert}\frac{\partial\ell_{n}(\widehat{\bm{\beta}})}{\partial\bm{\beta}_{g}}\bigg{\rVert}_{2}\geq\lambda_{\theta}^{\star}(\bm{0}_{m_{g}})\right)\rightarrow 0.$
@@
$P\left(\exists~g\in S_{0}^{c},~~\lVert h_{g}(\widehat{\bm{\beta}})\rVert_{2}\geq\lambda_{\theta}^{\star}(\bm{0}_{m_{g}})\right)$
$\qquad\leq P\left(\exists~g\in S_{0}^{c},~~\lVert h_{g}(\bm{\beta}_{0})\rVert_{2}\geq\frac{\lambda_{\theta}^{\star}(\bm{0}_{m_{g}})}{2}\right)+P\left(\exists~g\in S_{0}^{c},~~\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\rVert_{2}\geq\frac{\lambda_{\theta}^{\star}(\bm{0}_{m_{g}})}{2}\right)$
$\qquad\leq\sum_{g=s_{0}+1}^{G}P\left(\lVert h_{g}(\bm{\beta}_{0})\rVert_{2}\geq\frac{\lambda_{\theta}^{\star}(\bm{0}_{m_{g}})}{2}\right)+\sum_{g=s_{0}+1}^{G}P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\rVert_{2}\geq\frac{\lambda_{\theta}^{\star}(\bm{0}_{m_{g}})}{2}\right)$
$\qquad\overset{\Delta}{=}T_{1}+T_{2}.$ (A.21)
@@
$\leq 2\exp\left(-\frac{\lambda_{0}^{2}/(4D^{2}(\log G)^{2}m_{\max})}{2[2nG+L\lambda_{0}/(D{\color[rgb]{0,0,0}\log G}\sqrt{m_{\max}})]}\right)$
$=2\exp\left(-\frac{\lambda_{0}^{2}}{16D^{2}(\log G)^{2}m_{\max}nG+8L\lambda_{0}D\log G\sqrt{m_{\max}}}\right)$ (A.23)
@@
$T_{1}<2\exp\left(\log G-\frac{\lambda_{0}^{2}}{16D^{2}(\log G)^{2}m_{\max}nG+8L\lambda_{0}D\log G\sqrt{m_{\max}}}\right)\rightarrow 0,$ (A.24)
@@
We next examine each of the summands in $T_{2}$ in (A). In light of (A) that $\lambda_{\theta}^{\star}(\bm{0}_{m_{g}})>\lambda_{0}/2$, we have that for each $g\in S_{0}^{c}$ and sufficiently large $n$,
$P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\lVert_{2}\geq\frac{\lambda_{\theta}^{\star}(\bm{0}_{m_{g}})}{2}\right)$ $<P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\lVert_{2}\geq\frac{\lambda_{0}}{4}\right)$
$\leq P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})\rVert_{2,\infty}\lVert\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}\geq\frac{\lambda_{0}}{4}\right)$
$\leq P\left(\lVert\mathbf{X}_{S_{0}}^{\top}\bm{\Sigma}(\widetilde{\bm{\beta}})\mathbf{X}_{g}\rVert_{2,\infty}\lVert\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}\geq\frac{\lambda_{0}}{4}\right)$
$\lesssim\frac{4n\epsilon_{n}}{\lambda_{0}},$ (A.25)
To arrive at (A), we used the fact that $\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})=\mathbf{X}_{S_{0}}^{\top}\bm{\Sigma}(\widetilde{\bm{\beta}})\mathbf{X}_{g}$ in the second line. In the last line, we used Assumption (A3)(iii) that for all $g\in\mathcal{S}_{0}^{c}$, $\lVert\mathbf{X}_{S_{0}}^{\top}\bm{\Sigma}(\widetilde{\bm{\beta}})\mathbf{X}_{g}\rVert_{2,\infty}=O(n)$, the fact that $\lVert\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}=O_{p}(\epsilon_{n})$, and Markov’s inequality. From (A) and the fact that $\lambda_{0}\asymp G^{c},c>2$, and $G\gg n$, we have as an upper bound for $T_{2}$ in (A),
$T_{2}\prec\frac{4Gn\epsilon_{n}}{\lambda_{0}}\prec\frac{Gn}{G^{c}}=\frac{n}{G^{c-1}}\rightarrow 0\textrm{ as }n\rightarrow\infty.$ (A.26)